OpenStudy (anonymous):

2x-[3x+(x-2)] Acc. Algebra 2. Please help. Simplifying expressions.

OpenStudy (anonymous):

2x-[3x+(x-2)] = 2x-(3x+x-2) = 2x-(4x-2) = 2x-4x+2 = -2x+2 or 2-2x Sorry for the confusing previous post.

OpenStudy (anonymous):

2x-[3x+(x+2)] <---the addition sign infront of the bracket applies to the terms within it. So if we eliminate the bracket... =2x - [3x + x + 2] =2x - [4x + 2] =-2x +2 =-2 (x-1)

OpenStudy (anonymous):

Any one of your answers are correct but most people put the term with the x variable first and factor it as a "final" solution unless the instructions say otherwise.

OpenStudy (anonymous):

2*x - (3*x + (x - 2)) 2*x - (3*x + x - 2) 2*x - 3*x - x + 2 -x - x + 2 -2 x + 2 or 2-2x